OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the rate in feet per second.

OpenStudy (anonymous):

You need an angle. You mean straight up?

OpenStudy (anonymous):

Its vertical motion

Parth (parthkohli):

When the ball is at the ground, the height is \(0\). Replace \(h = 0\). \( \color{Black}{\Rightarrow 0 = 96t - 16t^2 }\)

OpenStudy (hba):

The ball hits the ground after 6 seconds.

OpenStudy (hba):

0 = 96ft/s * t - 16 t^2 16t^2 = 96t 16t = 96 t = 96/16 = 6

Parth (parthkohli):

\(0 = 16t(6 - t)\) \(0 = 6 - t\) Solve the rest.

OpenStudy (anonymous):

@ParthKohli how do i solve the rest?

Parth (parthkohli):

• Subtract 6 from both sides. • Multiply both sides by -1.

OpenStudy (anonymous):

so the answer is 6?

Parth (parthkohli):

Yeah!

OpenStudy (anonymous):

okay thanks!

Parth (parthkohli):

You're welcome, Megan O:)